OpenStudy (anonymous):

i could use some help please! translate the phrase into a mathematical expression and then simplify. five times a number decreased by twice the same number (type an expression using x as a variable. Do not simplify) the simplified expression is___

OpenStudy (anonymous):

5x - 2x = 3x 5(5) - 2(5) = 3(5) 25-10=15

OpenStudy (anonymous):

so what would the answer be?